SINGLE SHOT ULTRAFAST ELECTRON DIFFRACTION

\underline{F.M. RUDAKOV}, J.D. CARDOZA, P.M. WEBER, Department of Chemistry, Brown University, 324 Brook St, Providence, RI 02912, USA;; D.H. DOWELL, J.F. SCHMERGE, J.B. HASTINGS, Stanford Linear Accelerator, 2575 Sand Hill Rd, Menlo Park, CA 94025, USA.

The time resolution of ultrafast electron diffraction experiment is primarily limited by the electron pulse duration. The radio frequency (rf)-guns are known to be capable of creating sub-picosecond pulses of relativistic electrons. We performed both experimental and computational studies off electron diffraction using the rf-gun at the Gun Test Facility located at the Stanford Linear Accelerator. The copper photocathode of the rf-gun is illuminated by the 1mm diameter beam from a quadrupled Nd:Glass laser beam at 263 nm, with a pulse width of 2 ps. The electrons scatter from a piece of 160 nm thick aluminum foil placed at 0.755 m from the photocathode, which subsequently get detected 4.7 m downstream. We observed the diffraction pattern obtained with a single electron pulse of 2 pC (10^{7}) electrons) and having a duration of 500 fs. Our simulations show that by reducing the laser pulse duration and the charge per pulse the time resolution of 100 fs is readily attainable.